# Xmodel-VLM: A Simple Baseline for Multimodal Vision Language Model

### v2

@@ 1 Introduction
Despite the encouraging advancements made in the realm of visual language models, the pursuit of a genuinely optimal harmony between performance and efficiency remains an active and ongoing challenge. To this end, we trained a 1B scale language model from scratch, and guided by the cross-modal alignment principle advocated by the LLaVA paradigm, we have delved deeply into various facets of model architecture and training, including the selection of image encoders, the design of image-text connectors, and the exploitation of diverse datasets, all with the aim of pushing the boundaries of what is achievable with a smaller-scale model.
@@ 5.2 Token Numbers
Surprisingly, as the token count decreases, although there is an overall downward trend in various evaluation metrics, the change is remarkably gradual.

### v3

@@ 1 Introduction
Despite the encouraging advancements made in the realm of visual language models, the pursuit of a genuinely optimal harmony between performance and efficiency remains an active and ongoing challenge. To this end, we trained a 1B scale language model (Xmodel-LM-1B [41]) from scratch, and guided by the cross-modal alignment principle advocated by the LLaVA paradigm, we have delved deeply into various facets of model architecture and training, including the selection of image encoders, the design of image-text connectors, and the exploitation of diverse datasets, all with the aim of pushing the boundaries of what is achievable with a smaller-scale model.
@@ 5.2 Token Numbers
| 18 | 54.4 | 44.7 | 32.1 | 83.3 | 1247.1 | 40.2 |
| 8 | 52.3 | 40.8 | 29.8 | 80.7 | 1120.5 | 41.2 |
| 4 | 51.2 | 43.2 | 29.2 | 80.1 | 1040.1 | 36.2 |
| 2 | 49.7 | 44.0 | 28.8 | 79.6 | 1006.6 | 33.0 |
| 1 | 49.4 | 44.7 | 27.5 | 79.2 | 1070.6 | 32.1 |
Surprisingly, as the token count decreases, although there is an overall downward trend in various evaluation metrics, the change is remarkably gradual. Especially when only a small number of visual tokens are used, such as 1 and 2, there is no significant difference in the evaluation results. Of course, this may be related to the fact that there is no fine-grained description of the image in the datasets, which is the limitation of the vision encoder in modality alignment.